# Automatic Organization of Neural Modules for Enhanced Collaboration in Neural Networks

### v2

@@ Appendix C The Connection with DEQ
Appendix D The Connection with OPTNET
@@
Appendix E The supplementary materials for the experiments
In this section, we outline the parameters with the complexity of 100 nodes used in our experiments for each data set as shown in Table 5. Furthermore, for other complexity on other tasks. as demonstrated in the previous section, the parameters for $NM$ regularization and approximation can be tuned.
In our experiments, for the regular function $norm$, we use the absolute value of the weights. Like $L1$ regularization, we select this function with the aim of expanding the neural module and focus on several critical edges in each iteration. For larger neural modules, such critical edges might be rarer in the process of $NM$ regularization.
@@ Appendix E The supplementary materials for the experiments
Table 5: Parameters in the Experimants
Codon Usage Gases Concentration Postural Transitions
learningRate 0.1 0.1 0.1
momentum 0.5 0.5 0.5
scaling rate 1 1 1
$NM$ regularization $\alpha$ 0.2 0.2 0.2
percentage of approximation 0.05 0.01 0.01
Appendix F Neural Module Is Only for Approximation to Solve the System of Equations
@@
Note that, we update all the edges in every iteration, which provides additional benefits such as enhanced robustness. For the absolute value of the edge’s weight lower than $\gamma$, it also chances to grow larger than $\gamma$ in the coming iteration. That means every edge in the graph has the opportunity to change, thereby retaining the dynamism of Neural Modules and exerting our model’s performance.
Appendix G The Proof of the Universal Neural Module Approximation Theorem
@@
In the forward process, the system of equations constructs implicit functions across all nodes within the neural module. Subsequently, any node $n_{i}$ can be transformed into an explicit function. According to the Universal Approximation Theorem suc105 suc106 suc107, which states that for any continuous function on a compact set, there exists a one-hidden-layer feed-forward network capable of approximating the function, our system satisfies this condition. This is because our array of equations involves an activation function, fulfilling the requirement of the theorem. Therefore, the proof is complete.

### v3

@@ Appendix D The Connection with DEQ
Appendix E The Connection with OPTNET
@@
Appendix F The supplementary materials for the experiments
In this section, we provide a detailed description of the experimental setup, which is presented in Table 5. Additionally, as evidenced in this section, the parameters within our framework are adjustable, allowing for optimization and fine-tuning to achieve the best performance.
In our experimental setup, for the regularization function $norm$, we employ the absolute value of the weights, akin to $L1$ regularization. This choice is strategic, as it helps to expand the neural module while also highlighting a few critical edges during each iteration. In the context of larger neural modules, these critical edges may become less frequent as the process of $NM$ regularization unfolds. This approach encourages the model to concentrate on the most significant connections, thereby improving the efficiency and performance of the neural module.
@@ Appendix F The supplementary materials for the experiments
Table 5: Experimental Environment
CPU Gen Intel(R) Core(TM) i9-12900H 2.90 GHz
Cores 12
Memory 32G
GPU NVidia GeForce RTX 3060
Graphics Memory 12G
Norm Function L1
Other Parameters Refer to Figure 15
F.1 The Optimization of the Neural Modules
In this section, we explore the process of optimizing the structure of Neural Modules using $NM$ regularization. As discussed earlier, the results obtained have shown that $NM$ regularization would yield better results. Here, we investigate the parameter-setting strategies for our $NM$ regularization.
(a) Codon Usage Dataset
(b) Gases Concentration Dataset
(c) Postural Transitions Dataset
Figure 15: Parameter setting for $NM$ regularization.
The images in Figure 15 illustrate the impact of different parameters on the error rates of our $NM$ regulation across various percentages of approximations(controlled by parameter $\gamma$) for the Codon Usage Dataset. We found that the performance can be optimized by fine-tuning the $NM$ regularization and the percentages of approximations(controlled by parameter $\gamma$). Higher approximation percentages and regularization parameters produce more Neural Modules, which help reduce overfitting problems and in turn, enhance performance as well as robustness. However, overly sparse NMs could introduce extra bias and degrade performance. After standardization, the $NM$ regularization parameter with 100 nodes performs best at $0.2$ and the approximation on top $0.05$ edges by controlling parameter $\gamma$ for the Codon Usage Dataset. However, more stable performance appears on the approximation of top $0.03$ edges.
In Figure 15, we also offer this on the Gases Concentration Dataset and the Postural Transitions Dataset. We reached a similar conclusion that performance can be well optimized by tuning both the $NM$ regularization and the parameters of approximation. We observe that for these two datasets, the higher parameter $\gamma$ and $NM$ regularization parameters result in better performance. $NM$ regularization helps break large neural modules into smaller, more manageable neural modules. This segmentation leads to improved performance as previously introduced. However. compared with the Codeon Usage Data, the approximation can not to too large. On average, for 100 nodes, the $NM$ regularization parameter usually performs best at $0.2$ with the approximation on top $0.01$ edges by controlling parameter $\gamma$.
F.2 Trade off Performance against Efficient on Traditional NN
In this section, we propose an additional experiment to compare our framework with widely used traditional Neural Networks(NNs). The aim of this experiment is to assess how well our method performs in comparison to traditional NNs, especially when these networks have an adequate number of nodes, and to evaluate the trade-offs regarding efficiency. To prevent overfitting, we employ a more complex dataset with thousands of features. We utilize a web graph, a page-page graph of verified Facebook sites. In this graph, nodes represent official Facebook pages, and the edges represent mutual likes between these sites. This dataset can also be found in the UCI Machine Learning Repository.
(a) Performance NM vs.NN
(b) Running Time NM vs.NN. Measured in seconds.
Figure 16: NM vs. NN
From Image A Figure 16, we observed that the performance of traditional Neural Networks(NNs) deteriorates at 4000 nodes due to overfitting, which is also the best result it achieves. In contrast, our NM framework yields a more than 40 percent improvement in accuracy at 500 nodes and a even better result at 1000 nodes. On the other hand, as depicted in Figure B Figure 16, the improvement is with the count of several hours on efficiency. However, both performance and running time can be controlled by parameters such as sparsity. This indicates that our method is well-suited for tasks that prioritize accuracy. Users can weigh the trade-off between employing our method to enhance accuracy and the associated efficiency costs.
F.3 The Effect of $NM$ Regularization
To analyze the effect of $NM$ regularization, consider the probability of the edge’s weight being lower than $\gamma$ denoted as $\theta$ in the current iteration. For a neural module with $q$ nodes, the probability of a new edge integrating into the neural module would be denoted as
$1-\theta^{q}$ (11)
. In the case of a small $q$ value, $NM$ regularization seeks to increase the probability. Conversely, for larger $q$ values, NM regularization works to decrease the probability. Thus, NM regularization endeavors to minimize the disparity in the probability of a new edge integrating into the neural modules with varying sizes. This effect is illustrated in the first image of Figure 4.
Figure 17: Image A shows the analysis of $NM$ Regularization and image B displays the time-consumption for different sizes of neural modules which indicates parallel computing’s advantage.
Image A in Figure 17 illustrates the analysis of $NM$ Regularization and how balanced Neural Modules are formed through this process. Image B in Figure 17 shows the time consumption associated with different sizes of neural modules. As previously mentioned, the neural modules generated are independent and can be calculated in parallel. For instance, if a neural module comprises 50 nodes, the time to solve the system of equations required would be reduced to just 10ms in our experiments. Thus, $NM$ regularization enables sufficient management of larger node counts.
Appendix G Neural Module Is Only for Approximation to Solve the System of Equations
@@
Note that, we update all the edges in every iteration, which provides additional benefits such as enhanced robustness. For the absolute value of the edge’s weight lower than $\gamma$, it also chance to grow larger than $\gamma$ in the coming iteration. That means every edge in the graph has the opportunity to change, thereby retaining the dynamism of Neural Modules and exerting our model’s performance.
Appendix H The Proof of the Universal Neural Module Approximation Theorem
@@
In the forward process, the system of equations constructs implicit functions across all nodes within the neural module. Subsequently, any node $n_{i}$ can be transformed into an explicit function. According to the Universal Approximation Theorem [14] [18], which states that for any continuous function on a compact set, there exists a one-hidden-layer feed-forward network capable of approximating the function, our system satisfies this condition. This is because our array of equations involves an activation function, fulfilling the requirement of the theorem. Therefore, the proof is complete.